I need help with part D.

20pts for turned in on time) 8. (12pts)Carbon Dioxide Emissions. Burning fuels in power plants or motor vehicles emits carbon dioxide (CO2), which contributes to global warming. Table below shows Co2 emissions p person from countries with populations of at least 20 million. Carbon Dioxide emissions (metric tons per person) Co2 Ethiopia CO2 0.1I 0.1 Brazil 0.1 Egypt 0.2 Algeria 0.2 Iraq Indonesia 1.6 Ukraine 1.8 SouthAfrica Tanzania 2.6 Poland 2.9 Spain Myanmar Bangladesh 0.3 Turkey Sudan Nigeria Pakistan Philippines 0.3 Korea, North 3.3 Korea,South 0.3 Thailand 0.3 Argentina 0.4 Mexico 0.8 China 0.9 Romania 3.3 Japan 3.6 Germany 3.7 Russia 3.9 SaudiArabia 4.2 Canada 10.8 13.8 18.4 1 Uzbekistan 4.2Australia 5.4 United States 19.6 1 Venezuela 1.1 Malaysia 1.3 Iran 1.4 France Vietnam Colombia Morocco (4pts)Make a frequency histogram to display CO2 emissions per person. use classes width of 2 metric tons/person and starting with 0 metric tons/person. (Hint, to produce a histogram, you will need to generate a frequency by the closses in the frequency table) For the histogram, a. table first, then draw bars continuously 0.0-1.9 2.0-3.9 4.0-5.9 6.0-7.9 8.0-9.9 10.0-11.9 12.0-13.9 14.0-15.9 | 0 16.0-17.9 18.0-19.9 b. (2pt)What is the shape of the histogram (symmetric, skewed right, skewed left)? C. (2pts)which countries seem to be outliers from the histogram?
Homework #1-Chapter1 & 2 (20pts for turned in on time) The 5-number summary for the data is 0.1-0.95-3.3-7.4-19.6 (You can verity if desired). Use 1.5 IQR rule to find the outliers of the data, if there are any. Are they consistent with your findings in c, by checking with histogram? 12pts) Table blow shows annual spending on health care for different countries, with a unit of 1,000. Country Dollars 0.7 0.8 Poland Slovakia Korea Argentina Czech Republic Hungary Portugal Spain Greece Finland Germany Norway 1.1 1.1 1.3 1.3 1.8 1.9 2.0 2.1 2.8 3.4
Answer #1

d)
IQR = Q3-Q1
Q3=7.4, Q1=0.95
IQR = 7.4-0.95= 6.45

first quartile – 1.5·IQR = 0.95-1.5*6.45= -8.725
third quartile + 1.5·IQR = 7.4+1.5*6.45= 17.075

there are two data points 18.4 and 19.6 that does not lie between -8.725 and
17.075
hence these two are outliers